Consider the reaction 2H3PO4→P2O5+3H2O Using the information in the following table, calculate the average rate of formation of P2O5 between 10.0 and 40.0 s. Time (s) 0 10.0 20.0 30.0 40.0 50.0 [P2O5] (M)0 2.30×10−3 5.30×10−3 7.10×10−3 8.30×10−3 8.90×10−3.


1
Expert's answer
2020-06-11T10:31:27-0400

Solution.

"v = \pm\frac{dC}{dt}"

For products, average rate is positive. So,

"v = \frac{0.0083-0.0023}{40-10} = 0.0002 \ M \times L^{-1} \times s^{-1}"

Answer:

"v = 0.0002 \ M \times L^{-1} \times s^{-1}"